OpenStudy (anonymous):

if you saved $2 on Jan 1, and $4 on Feb 1 and $6 on Mar 1 and so on for a year how much would you save in a year?

OpenStudy (amistre64):

0$ if you account for the holidays

OpenStudy (amistre64):

2,4,6,8,10,12,14,16,18,20,22,24 12(2+24)/2 = 6(26) = 6(20+6) = 120 + 36 = 156 if i see it right